OpenStudy (anonymous):

Consider the motion of a projectile. It is fired at t = 0. Its initial speed is 11 m/s and its initial projection angle is 50◦ from the horizontal. How do I find the max height?

OpenStudy (amistre64):

11 sin(50) is your initial up velocity

OpenStudy (amistre64):

and since its in meters id go with g=4.9

OpenStudy (anonymous):

My homework tells me that g=9.8ms^2 why would you half that?

OpenStudy (amistre64):

-g/2 t^2 +Vi t + Hi where the highest point happens at: -Vi/-g

OpenStudy (amistre64):

its a math thing; they give you acceleration and you have to integrate it up to position

OpenStudy (anonymous):

Oh okay, I'll trust you on that bit. So, now what numbers go where? I found V, so, am I solving for time/ Or the variable Hi?

OpenStudy (amistre64):

a(t) = some constant: int a(t) = v(t) = c1 t + c2 int v(t) = h(t) = c*t^2/2 + c2 t + c3

OpenStudy (amistre64):

the Hi is initial height; lets assume 0 unless they give you it otherwise

OpenStudy (anonymous):

Its 0.

OpenStudy (amistre64):

Vi is the speed times the up vector of an angle of 50 degrees; up is sin 11 sin(50) t then

OpenStudy (amistre64):

and gravity ups to -, to get the thing back down to earth again, 9.8/2 t^2 to put it all toghter height(t) = -4.9 t^2 +11 sin(50)

OpenStudy (anonymous):

Why are we solving for initial velocity if the problem gave it to me as 11m/s?

OpenStudy (amistre64):

it didnt give us initial velocity perse; it gaves us the initial speed of the object going in a certain direction; at a 50 degree angle to the ground; that angle has 2 parts; out and up

OpenStudy (amistre64):

to find out how far up it goes, we need to isolate the up portion

OpenStudy (amistre64):



OpenStudy (amistre64):

up velocity = 11 sin(50) and you need to find out if that needs to be in radians or degrees

OpenStudy (amistre64):